OpenStudy (anonymous):

Which of the following statements have the same result? Explain each step in solving each one. f(1) when f(x) = 5x + 1 f−1(3) when f(x) = 2x plus 3, all over 5 3y − 7 = y + 5

OpenStudy (trojanpoem):

Use the equation button.

OpenStudy (trojanpoem):

\[f(1) , f(x) = 5x+1\] \[f^{-1}(3), f(x) = \frac{ 2x + 3 }{ 5 }\]

OpenStudy (trojanpoem):

f^-1(x) = (5f(x) - 3) / 2 = x x = f^-1(x) , f(x) = x f^-1(x) = (5x - 3)/2 f^-1(3) = 15-3/2 = 12/2 = 6 f(1) = 5 * 1 + 1 = 6 3y - 7 = y + 5 2y = 12 y = 6

OpenStudy (anonymous):

so 2 and 3 have the same answer. Thank you so much :)

OpenStudy (trojanpoem):

1, 2, 3 ?

OpenStudy (anonymous):

Oh. All 3 of them have the answer of 6?